OpenStudy (linda):

solve the equation 8x^3+4x^2-18x-9=0 algebraically for all values of x. help pleaseeee!!

OpenStudy (anonymous):

factor 4x^2 out of first two terms: 4x^2*(2x+1) factor (-9) out or other two terms: (-9)*(2x+1) Factor (2x+1) out of both terms you get: (2x+1)*(4x-9)=0 so, your solution - when one or other multp = 0, so 2x+1=0 and /or 4x-9=0 should be easy to solve it from this point. Let me know if there is a problem. Good Luck!

OpenStudy (radar):

When you factor out the (2x+1) you actually get: \[(4x ^{2}-9)(2x+1)\] which when further factored get: (2x+3)(2x-3)(2x+1) Now you have all three roots.

OpenStudy (linda):

ok thanks!